a mixture of 250cc of c2h6(150cc) and c2h4(100cc) was ignited with 500cc of oxygen .After a complete reaction,what will the composition of resulting mixture at 200degree c..
please solve at ur earliest

Grade:12th pass

1 Answers

deepak kumar
5 years ago
 the volume of c2h6 is 150ml and the  volume of c2h4 will is 100ml
now wrtinng reaction for their combustion
c2h6+7/2o2=2co2+3h2o
c2h4+3o2=2co2+2h2o
from reaction its clear that  volume of o2 used for combustion of c2h4 =100*3=300ml
remaining amount of o2=500-300=200ml
it will be used for combustion of c2h6
volume of c2h6 reacting =2/7*300=85.71ml
amount left of c2h6 =150-85.71=64.29ml
volume of co2 formed =10082+85.71*2=371.42
volume of water formed =100*2+85.71*3=457.13ml